# Achieving designed texture and flows in bulk active nematics using optimal control theory

### v1

@@ Appendix A Direct-adjoint-looping (DAL) method
We use direct-adjoint-looping (DAL), an iterative optimization method [101] to solve for the optimal schedule of activity in space and time that accomplishes our control goals. We start by writing the Lagrangian $\mathcal{L}$ of optimization, Eq. (6), where $\bm{\nu}$ and $\bm{\psi}$ act as Lagrange multipliers or adjoint variables that constrain the dynamics to follow Eqs. (1) and (2).
We construct an initial condition by performing a simulation with unperturbed dynamics (Eqs. (1) and (2)) until reaching steady-state, at a parameter set that leads to a desired initial behavior. Since our target states does not correspond to a steady state of the system, we define the target configuration by explicitly specifying the desired state variables. We also specify a time duration $t_{\text{F}}$ over which the control protocol will be employed, and an initial trial control protocol $\alpha^{0}(\mathbf{r},t)$. We then perform a series of DAL iterations, with each iteration involving the following steps:
@@
Iterations are continued until the gradient $\delta\mathcal{J}/{\delta\alpha}$ falls below a user-defined tolerance, which we set to $10^{-4}$. We employ Armijo backtracking [83] to adaptively choose the step-size $\Delta$ for gradient descent and to ensure convergence of the DAL algorithm.
@@ Appendix B Controllability
$\partial_{t}\delta\tilde{Q}_{xx}=-(iu_{x}^{0}q_{x}+iu_{y}^{0}q_{y})\delta\tilde{Q}_{xx}-Q_{xy}^{0}(iq_{y}\delta\tilde{u}_{x}-iq_{x}\delta\tilde{u}_{y})+\lambda iq_{x}\delta\tilde{u}_{x}+(a_{2}-a_{4}s)\delta\tilde{Q}_{xx}-Kq^{2}\delta\tilde{Q}_{xx}$ (12)
$\partial_{t}\delta\tilde{Q}_{xy}=-(iu_{x}^{0}q_{x}+iu_{y}^{0}q_{y})\delta\tilde{Q}_{xy}+Q_{xx}^{0}(iq_{y}\delta\tilde{u}_{x}-iq_{x}\tilde{u}_{y})+\frac{\lambda}{2}(iq_{x}\delta\tilde{u}_{y}+iq_{y}\delta\tilde{u}_{x})+(a_{2}-a_{4}s)\delta\tilde{Q}_{xy}-Kq^{2}\delta\tilde{Q}_{xy}$ (13)
@@
$\partial_{t}\begin{pmatrix}\delta\tilde{Q}_{xx}\\ \delta\tilde{Q}_{xy}\end{pmatrix}=$ $\begin{pmatrix}\lambda\frac{q_{x}^{4}}{q^{2}(q^{2}+\xi)}2\alpha+a_{2}-a_{4}s-Kq^{2}&\lambda\alpha\frac{q_{x}q_{y}}{q^{2}+\xi}(\frac{2q_{x}^{2}}{q^{2}}+1)\\ \frac{q_{x}q_{y}}{(q^{2}+\xi)}\alpha+\alpha\lambda\frac{q_{x}^{2}-q_{y}^{2}}{q^{2}}\frac{q_{x}q_{y}}{q^{2}+\xi}&-\frac{(q_{x}^{2}-q_{y}^{2})}{2(q^{2}+\xi)}\alpha+\frac{\alpha\lambda}{2(q^{2}+\xi)}(\frac{4q_{x}^{2}q_{y}^{2}}{q^{2}}+q^{2})+a_{2}-a_{4}s-Kq^{2}\end{pmatrix}$ (14)
@@
$A_{\text{q}}B_{\text{q}}=\left(\begin{array}[]{c}2\alpha\lambda^{2}\frac{q_{x}^{8}}{q^{4}(q^{2}+\xi)^{2}}+(a_{2}-a_{4}-Kq^{2})\lambda\frac{q_{x}^{4}}{q^{2}(q^{2}+\xi)}\\ +\frac{\lambda\alpha}{2}\frac{q_{x}^{2}q_{y}^{2}}{(q^{2}+\xi)^{2}}\left(\frac{2q_{x}^{2}}{q^{2}}+1\right)\left(1+\lambda\frac{q_{x}^{2}-q_{y}^{2}}{q^{2}}\right)\par\\[15.0pt] \left(\frac{q_{x}q_{y}}{q^{2}+\xi}\alpha+\alpha\lambda\frac{q_{x}^{2}-q_{y}^{2}}{q^{2}}\frac{q_{x}q_{y}}{q^{2}+\xi}\right)\lambda\frac{q_{x}^{4}}{q^{2}(q^{2}+\xi)}\\ +\left(\frac{\alpha\lambda}{2(q^{2}+\xi)}\left(\frac{4q_{x}^{2}q_{y}^{2}}{q^{2}}+q^{2}\right)-\frac{q_{x}^{2}-q_{y}^{2}}{2(q^{2}+\xi)}\right)\left(1+\lambda\frac{q_{x}^{2}-q_{y}^{2}}{q^{2}}\right)\frac{q_{x}q_{y}}{2(q^{2}+\xi)}\\ +(a_{2}-a_{4}s-Kq^{2})\left(1+\lambda\frac{q_{x}^{2}-q_{y}^{2}}{q^{2}}\right)\frac{q_{x}q_{y}}{2(q^{2}+\xi)}\par\par\end{array}\right).$
@@
$C=\begin{pmatrix}\lambda\frac{q_{x}^{2}}{q_{x}^{2}+\xi}&2\alpha\lambda^{2}\frac{q_{x}^{4}}{(q_{x}^{2}+\xi)^{2}}\\ &+\lambda(a_{2}-a_{4}s-Kq^{2})\frac{q_{x}^{2}}{q_{x}^{2}+\xi}\\[10.0pt] 0&0\end{pmatrix}$ (15)
The column vectors of this matrix are not linearly independent, hence the linear theory becomes uncontrollable when gradients perpendicular to direction of broken symmetry are absent. Thus, we see that spatial gradients orthogonal to the direction of the local alignment are critical to obtaining control solutions for an active nematic fluid.
@@
Figure 7: Coherent flow through a ‘channel’ without confining boundaries. The target state has uniform flow along the $\hat{x}$ direction with magnitude $U_{\text{T}}=0.5$ within the strip enclosed by green dashed lines. (a) The velocity field (arrows) and magnitude $u$ (color map) are shown at indicated time points for a representative trajectory under the optimal control protocol. (b) The activity field $\alpha$ (color map) and nematic director field (lines) are shown at the same time points. The initial state has the nematic uniformly aligned along the $\pi/4$ axis. (c) The flow speed $u_{x}$ (c) and the activity (d) averaged over the channel and bulk regions as a function of time. The parameter values for this solution were $\{J_{1},J_{2},J_{3},\zeta,W_{\text{u}},W_{\text{Q}},\alpha_{0},t_{\text{F}}\}=\{1.0,1.0,0.1,0.1,30.0,0,0.3,500.0\}$.
Appendix D Residues: deviations from target state as a function of time
@@
Appendix E Movie descriptions
  • Movie S1: Aligning the director field in a square subdomain (Fig. 3 main text). The left panel shows the nematic order parameter $s$ (color map) and director field $\mathbf{Q}$ (lines). The right panel shows the activity field $\omega$, with a logarithmic scale colorbar.
  • Movie S2: Coherent flow through a channel without confining boundaries (Fig. 4 main text). The top panel shows the x-velocity $\mathbf{u}_{x}$ (color map) and director field $\mathbf{Q}$ (lines). The bottom panel shows the activity field $\omega$, with a linear scale colorbar.
  • Movie S3: Couette flow in a channel without a moving boundary (Fig. 5 main text). The top panel shows the x-velocity $\mathbf{u}_{x}$ (color map) and director field $\mathbf{Q}$ (lines). The bottom panel shows the activity field $\omega$, with a linear scale colorbar.
  • Movie S4: Coherent flow through a channel without confining boundaries with high activity baseline ( Fig. 7). The top panel shows the x-velocity $\mathbf{u}_{x}$ (color map) and director field $\mathbf{Q}$ (lines). The bottom panel shows the activity field $\omega$, with a linear scale colorbar.

### v2

@@ Appendix A Direct-adjoint-looping (DAL) method
We use direct-adjoint-looping (DAL), an iterative optimization method Kenway et al. (2019) to solve for the optimal schedule of activity in space and time that accomplishes our control goals. We start by writing the Lagrangian $\mathcal{L}$ of optimization, Eq. (6), where $\bm{\nu}$ and $\bm{\psi}$ act as Lagrange multipliers or adjoint variables that constrain the dynamics to follow Eqs. (1) and (2).
We construct an initial condition by performing a simulation with unperturbed dynamics (Eqs. (1) and (2)) until reaching steady-state, at a parameter set that leads to a desired initial behavior. Since our target states are not steady states, we define the target configuration by explicitly specifying the desired state variables. We also specify a time duration $t_{\text{F}}$ over which the control protocol will be employed, and an initial trial control protocol $\alpha^{0}(\mathbf{r},t)$. We then perform a series of DAL iterations, with each iteration involving the following steps:
@@
Iterations are continued until the gradient $\delta\mathcal{J}/{\delta\alpha}$ falls below a user-defined tolerance, which we set to $10^{-4}$. We employ Armijo backtracking Borzì and Schulz (2011) to adaptively choose the step-size $\Delta$ for gradient descent and to ensure convergence of the DAL algorithm.
@@ Appendix B Controllability
$\partial_{t}\delta\tilde{Q}_{xx}=-(iu_{x}^{0}q_{x}+iu_{y}^{0}q_{y})\delta\tilde{Q}_{xx}-Q_{xy}^{0}(iq_{y}\delta\tilde{u}_{x}-iq_{x}\delta\tilde{u}_{y})+\lambda iq_{x}\delta\tilde{u}_{x}+(a_{2}-a_{4}s^{2})\delta\tilde{Q}_{xx}-Kq^{2}\delta\tilde{Q}_{xx}$ (13)
$\partial_{t}\delta\tilde{Q}_{xy}=-(iu_{x}^{0}q_{x}+iu_{y}^{0}q_{y})\delta\tilde{Q}_{xy}+Q_{xx}^{0}(iq_{y}\delta\tilde{u}_{x}-iq_{x}\tilde{u}_{y})+\frac{\lambda}{2}(iq_{x}\delta\tilde{u}_{y}+iq_{y}\delta\tilde{u}_{x})+(a_{2}-a_{4}s^{2})\delta\tilde{Q}_{xy}-Kq^{2}\delta\tilde{Q}_{xy}$ (14)
@@
$\partial_{t}\begin{pmatrix}\delta\tilde{Q}_{xx}\\ \delta\tilde{Q}_{xy}\end{pmatrix}=$ $\begin{pmatrix}\lambda\frac{q_{x}^{4}}{q^{2}(q^{2}+\xi)}2\alpha+a_{2}-a_{4}s^{2}-Kq^{2}&\lambda\alpha\frac{q_{x}q_{y}}{q^{2}+\xi}(\frac{2q_{x}^{2}}{q^{2}}+1)\\ \frac{q_{x}q_{y}}{(q^{2}+\xi)}\alpha+\alpha\lambda\frac{q_{x}^{2}-q_{y}^{2}}{q^{2}}\frac{q_{x}q_{y}}{q^{2}+\xi}&-\frac{(q_{x}^{2}-q_{y}^{2})}{2(q^{2}+\xi)}\alpha+\frac{\alpha\lambda}{2(q^{2}+\xi)}(\frac{4q_{x}^{2}q_{y}^{2}}{q^{2}}+q^{2})+a_{2}-a_{4}s^{2}-Kq^{2}\end{pmatrix}$ (15)
@@
$A_{\text{q}}B_{\text{q}}=\left(\begin{array}[]{c}2\alpha\lambda^{2}\frac{q_{x}^{8}}{q^{4}(q^{2}+\xi)^{2}}+(a_{2}-a_{4}s^{2}-Kq^{2})\lambda\frac{q_{x}^{4}}{q^{2}(q^{2}+\xi)}\\ +\frac{\lambda\alpha}{2}\frac{q_{x}^{2}q_{y}^{2}}{(q^{2}+\xi)^{2}}\left(\frac{2q_{x}^{2}}{q^{2}}+1\right)\left(1+\lambda\frac{q_{x}^{2}-q_{y}^{2}}{q^{2}}\right)\par\\[15.0pt] \left(\frac{q_{x}q_{y}}{q^{2}+\xi}\alpha+\alpha\lambda\frac{q_{x}^{2}-q_{y}^{2}}{q^{2}}\frac{q_{x}q_{y}}{q^{2}+\xi}\right)\lambda\frac{q_{x}^{4}}{q^{2}(q^{2}+\xi)}\\ +\left(\frac{\alpha\lambda}{2(q^{2}+\xi)}\left(\frac{4q_{x}^{2}q_{y}^{2}}{q^{2}}+q^{2}\right)-\frac{q_{x}^{2}-q_{y}^{2}}{2(q^{2}+\xi)}\right)\left(1+\lambda\frac{q_{x}^{2}-q_{y}^{2}}{q^{2}}\right)\frac{q_{x}q_{y}}{2(q^{2}+\xi)}\\ +(a_{2}-a_{4}s^{2}-Kq^{2})\left(1+\lambda\frac{q_{x}^{2}-q_{y}^{2}}{q^{2}}\right)\frac{q_{x}q_{y}}{2(q^{2}+\xi)}\par\par\end{array}\right).$
@@
$C=\begin{pmatrix}\lambda\frac{q_{x}^{2}}{q_{x}^{2}+\xi}&2\alpha\lambda^{2}\frac{q_{x}^{4}}{(q_{x}^{2}+\xi)^{2}}\\ &+\lambda(a_{2}-a_{4}s^{2}-Kq^{2})\frac{q_{x}^{2}}{q_{x}^{2}+\xi}\\[10.0pt] 0&0\end{pmatrix}$ (16)
The column vectors of this matrix are not linearly independent, hence the linear theory becomes uncontrollable when gradients perpendicular to the direction of broken symmetry vanish. Thus, we see that spatial gradients orthogonal to the direction of the local alignment are critical to obtaining control solutions for an active nematic fluid.
@@
Figure 9: Coherent flow through a ‘channel’ without confining boundaries. The target state has uniform flow along the $\hat{x}$ direction with magnitude $U_{\text{T}}=0.5$ within the strip enclosed by green dashed lines. (a) The velocity field (arrows) and magnitude $u$ (color map) are shown at indicated time points for a representative trajectory under the optimal control protocol. (b) The activity field $\alpha$ (color map) and nematic director field (lines) are shown at the same time points. The initial state has the nematic uniformly aligned along the $\pi/4$ axis. (c) The flow speed $u_{x}$ (c) and the activity (d) averaged over the channel and bulk regions as a function of time. The parameter values for this solution were $\{J_{1},J_{2},J_{3},\zeta,W_{\text{u}},W_{\text{Q}},\alpha_{0},t_{\text{F}}\}=\{1.0,1.0,0.1,0.1,30.0,0,0.3,500.0\}$. A video of this trajectory is provided in supplemental movie S5 SIr .
@@
Appendix D Movie descriptions
  • Movie S1: Aligning the director field in a square subdomain (Fig. 3 main text). The left panel shows the nematic order parameter $s$ (color map) and director field $\mathbf{Q}$ (lines). The right panel shows the activity field $\alpha$, with a logarithmic scale colorbar.
  • Movie S2: Coherent flow through a channel without confining boundaries (Fig. 4 main text) for $L_{y}=50$. The top panel shows the x-velocity $\mathbf{u}_{\mathbf{x}}$ (color map) and director field $\mathbf{Q}$ (lines). The bottom panel shows the activity field $\alpha$, with a linear scale colorbar.
  • Movie S3: Coherent flow through a channel without confining boundaries (Fig. 4 main text) for $L_{y}=100$. The top panel shows the x-velocity $\mathbf{u}_{\mathbf{x}}$ (color map) and director field $\mathbf{Q}$ (lines). The bottom panel shows the activity field $\alpha$, with a linear scale colorbar.
  • Movie S4: Couette flow in a channel without a moving boundary (Fig. 5 main text). The top panel shows the x-velocity $\mathbf{u}_{\mathbf{x}}$ (color map) and director field $\mathbf{Q}$ (lines). The bottom panel shows the activity field $\alpha$, with a linear scale colorbar.
  • Movie S5: Coherent flow through a channel without confining boundaries with high activity baseline (Fig. 9). The top panel shows the x-velocity $\mathbf{u}_{\mathbf{x}}$ (color map) and director field $\mathbf{Q}$ (lines). The bottom panel shows the activity field $\alpha$, with a linear scale colorbar.